OpenStudy (anonymous):

What is the value of x? 4x - 9 = 27 x =

OpenStudy (anonymous):

X=9

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

yeah his right

OpenStudy (anonymous):

or she

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

can i ask another?

OpenStudy (anonymous):

duhhh lol

OpenStudy (anonymous):

one sec lol

OpenStudy (anonymous):

just msg it to me

OpenStudy (anonymous):

What is the value of x? 6x - 2 (x +4) =12 x=

OpenStudy (anonymous):

x = 5 on that last one I think,

OpenStudy (anonymous):

@suppa_lova

OpenStudy (anonymous):

that would be 21 if it was 5

OpenStudy (anonymous):

X=1 because when you simplify it comes out as 6x-(2x+8)=12. then you subtract the 8 from both sides to get 6x-2x=4. then finally you combine like terms and get 4x=4 then simplify to get x=1

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

1 more What is the value of x? 8x-9=3x+16 x=

OpenStudy (anonymous):

8x-9=3x+16 is where you combine like terms to get 5x=25 so x=5

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

thank u so much i have a few more so mabye ill open another question!

OpenStudy (anonymous):

just msg them to me

OpenStudy (mrnood):

@suppa_lova 's answer above is WRONG "X=1 because when you simplify it comes out as 6x-(2x+8)=12. then you subtract the 8 from both sides to get 6x-2x=4. then finally you combine like terms and get 4x=4 then simplify to get x=1" He has not correctly distributed the -2 into the brackets it SHOULD be 6x-(2x+8)=12 6x-2x-8 =12 then ADD 8 to each side....